Pedro and Maria,10 miles a part,were walking towards each other.If the rate of PEDRO is 9mph and MARIA is 6mph,in how many minutes will they meet.?

Answer by jim_thompson5910(35256)

9x + 6x = 10

15x = 10

x = 10/15

x = 2/3 hours = 40 minutes


So it will take them 40 minutes til they meet.